40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 2 17:52:16 2011, kiwamu, Update, SUS, ITMX realigned 
    Reply  Tue Sep 6 08:05:53 2011, steve, Update, SUS, ITMX must be touching osemITMX.jpg
Message ID: 5341     Entry time: Tue Sep 6 08:05:53 2011     In reply to: 5337
Author: steve 
Type: Update 
Category: SUS 
Subject: ITMX must be touching 

Quote:

The new ITMX was aligned by changing the DC biases.

The resultant DC biases are reasonably small.

C1:SUS-ITMX_PIT_COMM = -0.2909

C1:SUS-ITMX_YAW_COMM = -0.0617

 

The alignment was done by trying to resonate the green light in the X arm cavity.

The spot position of the green light on the ITMX mirror looked good. This was confirmed by inserting a sensor card.

I did the OSEM mid-range adjustment and the rotation adjustment but at the end the OSEM DC voltage has changed due to the DC bias operation.

The OSEM rotation was approximately optimized so that all the face shadow sensors are sensitive to the POS motion but the SIDE shadow sensor is insensitive to the POS motion.

It needs a free swinging diagnosis.

 ITMX OSEMs   UL 1.8V,   UR 1.7V,   LR 0V,   LL 0V,   SD 1.3V  at the same bias setting shown above. May be a lose earth quake tip?or magnet is touching?

Attachment 1: osemITMX.jpg  98 kB  | Hide | Hide all
osemITMX.jpg
ELOG V3.1.3-